# Nature of phase transitions and metastability in scalar-tensor theories

### v1

@@ Abstract
$S_{\text{m}}$ is the matter action and $f_{\text{m}}$ denotes the matter fields which couple to the conformally scaled Jordan-frame metric. We use $A(\phi)=e^{\beta\phi^{2}/2}$ as in the original discovery paper of Damour and Esposito-Farèse [8], and add the scalar mass term $m_{\phi}$ [24]. Hence, our STTs live on a $(\beta,m_{\phi})$ parameter space. The scalar field obeys
$\Box\phi=\left(-4\pi\beta A^{4}\tilde{T}+m_{\phi}^{2}\right)\phi.$ (2)
The term in the parentheses becomes negative at high matter densities for $\beta<0$, which leads to a mode with purely imaginary oscillation frequency, a tachyon. Scalar fields grow exponentially around massive NSs due to this instability, which is eventually suppressed by nonlinear effects, leading to a star inside a stable scalar cloud, called a scalarized star. Note that an unscalarized star with $\phi=0$, a GR solution, is always an equilibrium solution for our choice of $A(\phi)$, but it is not stable if the tachyon is present.
@@
Let us start with the “standard” scalarization picture. The unscalarized solutions, which are also solutions in GR, are on the black curve, and the scalarized ones are on the turquoise one in Fig. 1. Scalarized solutions branch-off at a critical baryon mass $M_{\text{crit}}$, below which there is only one solution for a given baryon mass, an unscalarized one. This solution is known to be stable, hence, can be observed in principle. For $M_{\text{b}}>M_{\text{crit}}$ though, there are two possible solutions, one of them being scalarized  [^1] There is a $\phi\to-\phi$ symmetry for our choice of $A(\phi)$, which means each scalarized solution in the figure represents two configurations with opposite scalar field parity. We will be talking about points on the scalarized branches of the $M_{\text{b}}(\tilde{\rho}_{\text{c}})$ curve as if they are single solutions to avoid confusion, since they look as such on the $M_{\text{b}}(\tilde{\rho}_{\text{c}})$ curves.. However, we have seen that the unscalarized solution is unstable to the growth of a scalar mode, hence it would not be observed. Thus, any existing star with these higher masses would be a scalarized one. We call this picture second-order scalarization. The naming will become clear in our phase transition discussion.
@@
When there are multiple equilibrium solutions that are robust against small perturbations, as is our case, we have the phenomenon of metastability. We typically call the solution with the globally lowest energy to be the stable one (or the global minimum), whereas the others are metastable. A metastable state would energetically prefer to transition to the global minimum, but this can only occur for large enough perturbations which may or may not exist in a star’s environment. Hence, both solutions are possible in reality, which makes novel astrophysical scenarios possible as we will discuss later.
@@
We closely investigated only two points on the $(\beta,m_{\phi})$ parameter space of our STT so far, but what is the prevalence of first- and second-order scalarization in general? Fig. 3 (left) shows the maximum ADM mass difference between two stars with the same baryon mass, $\Delta M$, which is a crude approximation for the energy that would be released in a transition between two states where there is no matter loss. Since it means there are metastable states, any point with a nonzero value means scalarization is first order if the STT with parameters $(\beta,m_{\phi})$ governs gravity. One can see that second-order scalarization is the dominant type even for the original scalarization model for which $m_{\phi}=0$ [8]. This clearly demonstrates our claim about the ubiquity of first-order scalarization.
@@
In terms of detectability, realtively higher NS masses and higher $\Delta M$ are both desirable. This means that the prominent regions in Fig. 3 are not necessarily interesting on their own, but rather the middle region where $\Delta M$ and $M_{\text{ADM,branch}}$ are both relatively high likely provides the best observational prospects. Consider these results in the context that they are only for a specific STT. $\Delta M$ and $M_{\text{ADM,branch}}$ would vary for other, more general, STTs, providing opportunities for testing deviations from GR.
Figure 3: Dependence of scalarization characteristics on the theory parameters $(\beta,m_{\phi})$ for first order scalarization. There is no scalarization at all above the red curve [25]. Second-order scalarization occurs only in the white regions below the curve. Thus, first-order scalarization is by far the more likely outcome if scalarization occurs. Left: $\Delta M$, the maximum ADM mass difference between two stars of equal baryon mass. Right: ADM mass at the scalarization branch-off point. Both are in units of $M_{\odot}$.
Phenomenological descriptions of scalarization: The onset of scalarization has been viewed as a phase transition since the early days [10], and phenomenological explanations in terms of Landau theory have been occasionally invoked in the literature [31, 33]. Nevertheless, this has been in the context of second-order phase transitions, which only holds for second-order scalarization. [^2] The end, rather than the beginning, of scalarization was called a first-order phase transition in Kuan et al. [18], though this study was not specifically concerned with a phenomenological analysis. Here, we show that all unusual features of first-order scalarization in the $M_{\text{bottom}}<M_{\text{b}}<M_{\text{crit}}$ region can be explained by a first-order phase transition.
@@
First-order scalarization occurs when $b(M_{\text{b}})<0$, for which $c(M_{\text{b}})>0$ is also relevant. The behavior of $M_{\text{ADM}}(Q)$ changes drastically depending on the parameter $\chi\equiv ac/b^{2}$ as seen in Fig. 4. To start with, for $\chi>1/4$, the only equilibrium point and the global minimum of Eq. (3) is at $Q=0$. There is only a single stable, unscalarized configuration for a given $M_{\text{b}}$, corresponding to the region to the left of the line $A$ in Fig. 2.
As $M_{\text{b}}$ increases, $\chi$ drops below $1/4$ and local maxima (--) and outer local minima ($+$) appear at
$Q^{2}_{\pm}=\left[-b/(2c)\right]\left(1\pm\sqrt{1-4\ ac/b^{2}}\ \right).$ (4)
However, the global minimum is still located at $Q=0$ if $\chi>3/16$. This is the case where the unscalarized ($Q=0$) solution is the globally stable one, and the stable scalarized solutions are metastable. In Fig. 2, this corresponds to the region between vertical lines A and B. The local maxima are also equilibrium solutions, they are the unstable scalarized configurations (dotted line in Fig. 2) which have the highest total energy, i.e., the lowest binding energy among the equilibria.
As $M_{\text{b}}$ increases further, we arrive at $\chi=3/16$, where the energies at $Q=0$ and the outer minima become equal. This corresponds to line B in Fig. 2, and it is the actual point of the first-order phase transition. If $\chi$ becomes less than $3/16$, i.e., to the right of line B, the scalarized solution at $Q_{+}$ becomes the globally stable one, and the unscalarized stars become metastable. Note that, at the point of transition, the global minimum jumps discontinuously from $Q=0$ to one of $Q=\pm Q_{+}$, which is a hallmark of first-order phase transitions.
Further increasing $M_{\text{b}}$ brings us to $a=0$, for which the maxima at $Q=\pm Q_{-}$ join the local minimum at $Q=0$. This corresponds to the line C in Fig. 2. For negative values of $a$, we have a local maximum at $Q=0$, and minima at $Q=\pm Q_{+}$. This corresponds to having one unstable GR solution and one stable scalarized solution, which is the region to the right of line C.
The phase transition picture elucidates many points in our numerical results which previously had no clear reason. For example, unstable solutions are not necessarily energetically disfavored in general, but the unstable scalarized stars always have the lowest binding energy in Fig. 2. In Fig. 4, this is a trivial point since a local maximum between two local minima has to have the highest energy. The list continues for other features like (i) the stable scalarized stars being metastable at lower $M_{\text{b}}$ and becoming the global minimum only with increasing $M_{\text{b}}$, (ii) the branch continuously connected to the GR solutions being unstable and (iii) the unstable scalarized stars always having lower scalar charge compared to the scalarized ones: $Q_{-}<Q_{+}$ (not plotted here).
@@
Our phase transition picture already has a simple and powerful explanation for the above results. Changing the $\phi^{4}$ term in the expansion of $A(\phi)$ directly modifies the higher order expansion coefficients in our Landau ansatz (3), specifically $b$. Recalling that the main difference between a first and second-order transition is the sign of $b$, we can see how these higher order terms can determine the order of the phase transition. However, we emphasize that explicitly modifying the coupling or some form of fine tuning is not necessary for first-order scalarization by any means, as is clear from our choice of $A(\phi)$ here. Overall, we aim to study first-order scalarization for the most general STTs in order to devise more comprehensive tests for these theories.
@@
In terms of observational prospects, the specific examples we gave are most relevant for lower mass NSs since we concentrated on the onset of scalarization. There are recently proposed theoretical scenarios for these such as formation of NSs down to masses of $0.1M_{\odot}$ in accretion disks [40], and tidal disruption events in black hole–NS mergers where part of the NS mass can be ripped out [41, 42]. How a first-order transition would behave during the birth of a NS in a supernova explosion is also not known in detail [11], which might be relevant. There is already some work about first-order phase transition near the maximum allowed NS masses, for which possible astrophysical signals have been studied [18]. We expect these studies to be readily adaptable to the transitions at lower stellar masses we studied above. Lastly, we reiterate that what we considered in action (1) is the original and most popular scalarization theory, but it is one scalar coupling choice for a specific $A(\phi)$. We expect other models of scalarization to have different characteristics, where different NS mass ranges are possible or even dominant for first-order scalarization. Studying these more general theories and how to test them via discontinuous phase transitions are major future research directions.
@@
The phase transition order depends on the sign of $b$. Let us remark that the parameters $a,b,c$, as well as the function $M_{0}$ and the critical mass $M_{\text{crit}}$ depend on the parameters of the theory, $(\beta,m_{\phi})$, and also the EOS. For simplicity let us focus on the $\beta$ dependence, i.e. $b=b(M_{\text{b}},\beta)$. The fact that the order of the phase transition changes at some point as $\beta$ becomes more and more negative means that $b$ changes sign as $\beta$ drops below a critical value $\beta_{\text{crit}}<0$. When $\beta=\beta_{\text{crit}}$, we have a point where $a$ and $b$ vanish simultaneously, i.e., the point $(\beta=\beta_{\text{crit}},M_{\text{b}}=M_{\text{crit}}(\beta_{\text{crit}}))$. Such a point is called a tricritical point [35, 36], and it separates a line of second-order phase transitions on the $(\beta,M_{\text{b}})$ plane from a line of first-order ones. Spontaneous scalarization at $\beta_{\text{crit}}$ features special characteristics whose study is another possible future research topic. Similarly, other aspects of the voluminous phase transition literature can be a source of other surprising results for gravitational phase transitions.
@@
This study was supported by the Scientific and Technological Research Council of Turkey (TÜBİTAK) Grant Number 122F097. Authors thank TÜBİTAK for its support, and Daniela Doneva and Stoytcho Yazadjiev for helpful discussions.

### v2

@@ Abstract
$S_{\text{m}}$ is the matter action and $f_{\text{m}}$ denotes the matter fields which couple to the conformally scaled Jordan-frame metric. We use $A(\phi)=e^{\beta\phi^{2}/2}$ as in the original discovery paper of Damour and Esposito-Farèse [8], and add the scalar mass term $m_{\phi}$ [24]. Hence, our STTs live on a $(\beta,m_{\phi})$ parameter space. Note that $m_{\phi}$ has the important consequence of suppressing the scalar dipole radiation [25, 24, 9], which means our model here is not affected by various observations that severely constrain the original massless scalarization idea [26, 27, 28].
The scalar field obeys
$\Box\phi=\left(-4\pi\beta A^{4}\tilde{T}+m_{\phi}^{2}\right)\phi,$ (2)
$\tilde{T}$ being the trace of the stress-energy tensor in the Jordan frame. The term in the parentheses becomes negative at high matter densities for $\beta<0$, which leads to a mode with purely imaginary oscillation frequency, a tachyon. Scalar fields grow exponentially around massive NSs due to this instability, which is eventually suppressed by nonlinear effects, leading to a star inside a stable scalar cloud, called a scalarized star. Note that an unscalarized star with $\phi=0$, a GR solution, is always an equilibrium solution for our choice of $A(\phi)$, but it is not stable if the tachyon is present.
@@
Let us start with the “standard” scalarization picture. The unscalarized solutions, which are also solutions in GR, are on the black curve, and the scalarized ones are on the turquoise one in Fig. 1. Scalarized solutions branch off at a critical baryon mass $M_{\text{crit}}$, below which there is only one solution for a given baryon mass; an unscalarized one. This solution is known to be stable, hence, can be observed in principle. For $M_{\text{b}}>M_{\text{crit}}$ though, there are two possible solutions, one of them being scalarized  [^1] There is a $\phi\to-\phi$ symmetry for our choice of $A(\phi)$, which means each scalarized solution in the figure represents two configurations with opposite scalar field parity. We will be talking about points on the scalarized branches of the $M_{\text{b}}(\tilde{\rho}_{\text{c}})$ curve as if they are single solutions to avoid confusion, since they look as such on the $M_{\text{b}}(\tilde{\rho}_{\text{c}})$ curves.. However, we have seen that the unscalarized solution is unstable to the growth of a scalar mode, hence it would not be observed. Thus, any existing star with these higher masses would be a scalarized one. We call this picture second-order scalarization. The naming will become clear in our phase transition discussion.
@@
When there are multiple equilibrium solutions that are robust against small perturbations, as is our case, we have the phenomenon of metastability. We typically call the solution with the globally lowest energy the stable one (or the global minimum), whereas the others are metastable. A metastable state would energetically prefer to transition to the global minimum, but this can only occur for large enough perturbations which may or may not exist in a star’s environment. Hence, both solutions are possible in reality, which makes novel astrophysical scenarios possible as we will discuss later.
@@
We closely investigated only two points on the $(\beta,m_{\phi})$ parameter space of our STT so far, but what is the prevalence of first- and second-order scalarization in general? Fig. 3 (left) shows the maximum ADM mass difference $\Delta M$ between two stars with the same baryon mass, which is a crude approximation for the energy that would be released in a transition between two states where there is no matter loss. Since it means there are metastable states, any point with a nonzero value means scalarization is first order if the STT with parameters $(\beta,m_{\phi})$ governs gravity. One can see that first-order scalarization is the dominant type even for the original scalarization model for which $m_{\phi}=0$ [8]. This clearly demonstrates our claim about the ubiquity of first-order scalarization.
@@
In terms of detectability, relatively higher NS masses and higher $\Delta M$ are both desirable. This means that the prominent regions in Fig. 3 are not necessarily interesting on their own, but rather the middle region where $\Delta M$ and $M_{\text{ADM,branch}}$ are both relatively high likely provides the best observational prospects. Consider these results in the context that they are only for a specific STT. $\Delta M$ and $M_{\text{ADM,branch}}$ would vary for other, more general, STTs, providing opportunities for testing deviations from GR.
Figure 3: Dependence of scalarization characteristics on the theory parameters $(\beta,m_{\phi})$ for first order scalarization. There is no scalarization at all above the red line [29]. Second-order scalarization occurs only in the white regions below the red line. Thus, first-order scalarization is by far the more likely outcome if scalarization occurs. Left: $\Delta M$, the maximum ADM mass difference between two stars of equal baryon mass. Right: ADM mass at the scalarization branch-off point. Both are in units of $M_{\odot}$. $\Delta M<10^{-5}M_{\odot}$ is indistinguishable from numerical noise, hence excluded from the plot. We use lower resolution away from the boundary between first- and second-order scalarization, where fine detail is not critical. See “End Matter” for details.
Phenomenological descriptions of scalarization: The onset of scalarization has been viewed as a phase transition since the early days [10], and phenomenological explanations in terms of Landau theory have been occasionally invoked in the literature [35, 37]. Nevertheless, this has been in the context of second-order phase transitions, which only holds for second-order scalarization  [^2] Muniz and Mendes [47] is an exception, which was posted after the submission of our manuscript. The end, rather than the beginning, of scalarization was called a first-order phase transition in Kuan et al. [18], though this study was not specifically concerned with a phenomenological analysis.. Here, we show that all unusual features of first-order scalarization in the $M_{\text{bottom}}<M_{\text{b}}<M_{\text{crit}}$ region can be explained by a first-order phase transition.
@@
First-order scalarization occurs when $b(M_{\text{b}}){<}0$, for which $c(M_{\text{b}}){>}0$ is also relevant. The behavior of $M_{\text{ADM}}(Q)$ changes drastically depending on the parameter $\chi\equiv ac/b^{2}$ that decreases with increasing $M_{\text{b}}$, as seen in Fig. 4. There is a single stable equilibrium at $Q=0$ when $\chi$ is positive and high, which becomes unstable for negative $\chi$ for which two symmetric stable equilibria exist. In between these phases, we have three local minima corresponding to stable or metastable states, as well as two symmetric local maxima that correspond to unstable equilibria. Quantitative details can be found in the “End Matter.”
The phase transition picture in Fig. 4 elucidates many points in our numerical results which previously had no clear reason. For one, the unstable scalarized stars always have the lowest binding energy in Fig. 2, even though there is no necessity for unstable equilibria to be energetically disfavored. This becomes a trivial point in Fig. 4 since a local maximum between two local minima has to have the highest energy. The list continues for other features like (i) the stable scalarized stars being metastable at lower $M_{\text{b}}$ and becoming the global minimum only with increasing $M_{\text{b}}$, (ii) the branch continuously connected to the GR solutions being unstable and (iii) the unstable scalarized stars always having lower scalar charge compared to the scalarized ones: $Q_{-}<Q_{+}$ (not plotted here).
The main factor controlling the order of the phase transition is the sign of the coefficient $b(M_{\text{b}})$ in the Landau ansatz (3). For various theory parameters, we extracted the value of $b$ at the point where the scalarized NSs branch off from the GR solutions, $b_{0}=b(M_{\text{crit}})$. See “End Matter” for details.
Fig. 5 clearly shows how $b_{0}$ becomes more negative as $\beta$ becomes more negative for $m_{\phi}=0$, the case being similar for all scalar masses. This is exactly in line with our observations in Figs. 3 and 4.
Figure 5: $b_{0}=b(M_{\text{crit}})$ as a function of $\beta$ for $m_{\phi}=0$. The change from first- to second-order scalarization is near $\beta=-10.6$, that is, $b_{0}(\beta=-10.6)=0$.
@@
Our phase transition picture already has a simple and powerful explanation for the above results. Changing the $\phi^{4}$ term in the expansion of $A(\phi)$ directly modifies the higher order expansion coefficients in our Landau ansatz (3), specifically $b$. Recalling that the main difference between a first- and second-order transition is the sign of $b$, we can see how these higher order terms can determine the order of the phase transition. However, we emphasize that explicitly modifying the coupling or some form of fine tuning is not necessary for first-order scalarization by any means, as is clear from our choice of $A(\phi)$ here. Overall, we aim to study first-order scalarization for the most general STTs in order to devise more comprehensive tests for these theories.
@@
In terms of observational prospects, the specific examples we gave are most relevant for lower mass NSs since we concentrated on the onset of scalarization. There are recently proposed theoretical scenarios for these such as formation of NSs down to masses of $0.1M_{\odot}$ in accretion disks [44], and tidal disruption events in black hole–NS mergers where part of the NS mass can be ripped out [45, 46]. Such a change in the stellar mass can radically affect how large a scalar field can be supported by a star in general, and this becomes especially relevant for first-order scalarization where the discontinuous nature of the scalarization/descalarization process can provide distinct signals. There is already some work about first-order phase transitions near the maximum allowed NS masses, for which possible astrophysical signals have been studied [18]. For instance, a sudden transition between scalarized and non-scalarized configurations results in the emission of gravitational waves which carry two additional modes beyond those in GR. Moreover, these can also persist for longer periods. Together with related electromagnetic and neutrino signals from such a violent event, these gravitational waves offer distinct observational indicators for a first-order transition. Due to its discontinuous nature, first-order scalarization might also feature novel phenomenology in stellar collapse events, which have only been studied for the continuous second-order scalarization so far [11].
We reiterate that action (1) is the original and most popular scalarization theory in terms of the scalar coupling function $A(\phi)$, but it is still one specific choice. We expect other models of scalarization to have diverse characteristics, where different NS mass ranges are possible or even dominant for first-order scalarization. Studying these more general theories and how to test them via discontinuous phase transitions are major future research directions. Also note that the mass term $m_{\phi}$ makes almost all of our parameter space, the region with $m_{\phi}\gtrsim 10^{-15}$eV, unconstrained with current observations [24, 9]. We included the mostly-constrained $m_{\phi}=0$ case in our analysis for the sake of completeness and also because it is still informative since the $m_{\phi}\to 0$ limit is continuous in terms of the NS structure.
@@
The phase transition order depends on the sign of $b$. Let us remark that the parameters $a,b,c$, as well as the function $M_{0}$ and the critical mass $M_{\text{crit}}$ depend on the parameters of the theory, $(\beta,m_{\phi})$, and also the EOS. For simplicity let us focus on the $\beta$ dependence, i.e. $b=b(M_{\text{b}},\beta)$. The fact that the order of the phase transition changes at some point as $\beta$ becomes more and more negative means that $b$ changes sign as $\beta$ drops below a critical value $\beta_{\text{crit}}<0$, see Fig. 5. When $\beta=\beta_{\text{crit}}$, we have a point where $a$ and $b$ vanish simultaneously, i.e., the point $(\beta=\beta_{\text{crit}},M_{\text{b}}=M_{\text{crit}}(\beta_{\text{crit}}))$. Such a point is called a tricritical point [39, 40], and it separates a line of second-order phase transitions on the $(\beta,M_{\text{b}})$ plane from a line of first-order ones. Spontaneous scalarization at $\beta_{\text{crit}}$ features special characteristics whose study is another possible future research topic. Similarly, other aspects of the voluminous phase transition literature can be a source of other surprising results for gravitational phase transitions.
Lastly, we note that even though a phenomenological fit with very few parameters is possible, there is no understanding for the behavior of $b_{0}$ in Fig. 5 from first principles to the best of our knowledge. Recall that the scalar field profiles depend on the nonlinear terms that suppress the tachyonic instability, hence there is no simple method to relate the energy of the spacetime and the related Landau ansatz parameters to the STT parameters. Qualitatively, first-order scalarization becomes dominant when scalarization itself is stronger in terms of how much deviation we have from GR, which happens at higher $|\beta|$ [24, 9]. Whether this can lead us to the underlying reason for the change of sign in $b_{0}$ remains to be seen.
@@
This study was supported by the Scientific and Technological Research Council of Turkey (TÜBİTAK) Grant Number 122F097. Authors thank TÜBİTAK for its support, and Daniela Doneva and Stoytcho Yazadjiev for helpful discussions. Computations were performed on the KUACC cluster of Koç University.